# Fair Minimum Labeling: Efficient Temporal Network Activations for Reachability and Equity

### v1

@@ Abstract
Balancing resource efficiency and fairness is critical in networked systems that support modern learning applications. We introduce the Fair Minimum Labeling (FML) problem: the task of designing a minimum-cost temporal edge activation plan that ensures each group of nodes in a network has sufficient access to a designated target set, according to specified coverage requirements. FML captures key trade-offs in systems where edge activations incur resource costs and equitable access is essential, such as distributed data collection, update dissemination in edge-cloud systems, and fair service restoration in critical infrastructure. We show that FML is NP-hard and $\Omega(\log|V|)$-hard to approximate, and we present probabilistic approximation algorithms that match this bound, achieving the best possible guarantee for the activation cost. We demonstrate the practical utility of FML in a fair multi-source data aggregation task for training a shared model. Empirical results show that FML enforces group-level fairness with substantially lower activation cost than baseline heuristics, underscoring its potential for building resource-efficient, equitable temporal reachability in learning-integrated networks.
@@ 1 Introduction
The introduction of FML as a novel optimization problem is motivated by both practical relevance and theoretical gaps. First, it applies to a wide range of systems where temporal resource allocation directly affects group equity. In distributed data gathering for learning applications, such as those involving environmental sensors, mobile agents, or geographically diverse data streams [28, 36, 35], FML enables minimal-cost activation plans that ensure all groups are fairly represented in the collected data. In edge-cloud systems, it helps decide when communication links should be active so that all device groups can receive timely model updates under bandwidth constraints. Similarly, in infrastructure restoration, FML can guide activation sequences that ensure fair reactivation of services. Second, FML fills a critical gap in the literature. Existing fairness-aware graph methods, such as fair clustering [23], node embeddings [30], and influence maximization [33, 37], are primarily designed for static settings and do not model temporal interactions or activation cost. Fairness has been studied in wireless systems [16], and temporal connectivity has been explored in dynamic networks [26], but fairness constraints remain absent in algorithmic temporal graph design. FML addresses this gap by explicitly minimizing temporal edge activations, treated as a proxy for resource cost, while enforcing group-level temporal reachability. In contrast to decentralized approaches such as classical federated learning [18], where participation is uncoordinated, FML assumes centralized control over link activations, allowing the enforcement of fairness guarantees over time. This design objective distinguishes FML from prior work focused on static structures or continuous flows.
@@ 2 Preliminaries
A static graph is an ordered pair $G=(V,E)$, where $V$ is a finite set of nodes (or vertices) and $E\subseteq\{\{u,v\}\mid u,v\in V,u\neq v\}$ is a finite set of undirected edges. We assume static graphs are simple (no self-loops, no multiple edges between the same pair of nodes). A colored static graph $G=(V,E,c)$ with colors $\mathcal{C}$ has an additional function $c:V\rightarrow\mathcal{P}(\mathcal{C})$ assigning a (possibly empty) set of colors $c(u)\in\mathcal{C}$ to each node $u\in V$ where $\mathcal{C}=[c_{\max}]$ is a set of $c_{\max}$ colors. We use $V_{c}=\{v\in V\mid c(v)=c\}$. Let $n=|V|$ and m=|E|m=|{E}|.
A temporal graph is an ordered pair $\mathcal{G}=(V,\mathcal{E})$, where $V$ is a finite set of nodes and $\mathcal{E}$ is a finite set of temporal edges. Each temporal edge is a triple $e=(\{u,v\},\tau)$, where $\{u,v\}$ is an undirected pair of distinct nodes from $V$, and $\tau\in\mathbb{N}$ is the timestamp indicating when the interaction between $u$ and $v$ occurs.
Given a static graph $G=(V,E)$, a temporal labeling is a function $\lambda:E\rightarrow\mathcal{P}(\mathbb{N})$ that assigns a set of timestamps (possibly empty) to each static edge $e\in E$. A labeling $\lambda$ induces a temporal graph $G_{\lambda}=(V,E_{\lambda})$, where the node set is the same as the static graph’s, and the set of temporal edges is defined as: $E_{\lambda}=\{(\{u,v\},\tau)\mid e=\{u,v\}\in E\text{ and }\tau\in\lambda(e)\}$. The size (or cost) of a labeling $\lambda$, denoted by $|\lambda|$, is the total number of assigned timestamps, which corresponds to the number of temporal edges in the induced temporal graph is $|\lambda|=|E_{\lambda}|=\sum_{e\in E}|\lambda(e)|$.
@@
By convention, we consider every node to be temporally reachable from itself (via a path of length zero), so $r(v,v)=1$ for all $v\in V$. The specific temporal graph $\mathcal{G}$ (often an induced graph $G_{\lambda}$) for which reachability is considered will always be clear from the context. Finally, let $r(U,W)=\sum_{u\in U,v\in W}r(u,v)$ be the the sum of pairwise reachability of nodes in $U$ to nodes in $W$.
@@ 3 Related Work
Temporal Graph Connectivity and Labeling. A core component of FML is finding a minimum-cost set of temporal edge activations. Mertzios et al. [26] pioneered the study of temporal network design, where the goal is to label static graph edges with one or more timestamps to achieve specific temporal reachability properties. The Minimum Labeling (ML) problem is a temporal network design problem which seeks the minimum number of such labels to make an entire static graph temporally connected. For its directed version on strongly connected digraphs, [26] had provided bounds on the number of labels needed, proving that 2 labels per arc and at most 4(|V|1)4(|V|-1) total labels are always sufficient. Using similar constructions on undirected graphs, Akrida et al. [2] proved that 2|V|32|V|-3 total labels are sufficient, later refined by Klobas et al. [24] (to 2|V|42|V|-4 under certain conditions), who also showed ML is polynomially solvable. While ML aims for global temporal connectivity, FML focuses on targeted reachability for specific node groups to a designated terminal set, under explicit fairness (coverage) constraints for each group, and a cost minimization objective. Klobas et al. [24] also introduced extensions like Minimum Steiner Labeling (MSL), where the goal is to pairwise temporally connect a predefined set of terminal nodes. Our FML generalizes ML and MSL. Other research has explored concepts like temporally-connected spanning subgraphs [3], and various notions of temporal paths and reachability [21, 7, 27, 15]. However, these works generally do not incorporate group fairness constraints or the specific cost model of minimizing temporal edge activations as FML does.
@@
Fairness in Graph Mining and Network Optimization. There is a significant body of work on incorporating fairness into static graph algorithms and network optimization, driven by the need for equitable outcomes in various applications. This includes, e.g., fair clustering [1, 10, 23], fair node embeddings [30, 6], fair influence maximization [33, 37], and graph learning tasks [9, 11, 29, 38]. While these contributions are crucial for advancing fairness in graph-based AI, they predominantly operate on static graphs and often define fairness differently (e.g., proportionality in output, balanced representation in a static structure). For example, a related problem in static graphs is the Balanced Connected Subgraph (BCS) problem [5] which seeks a connected subgraph with balanced representation from different node colors. FML, in contrast, addresses fairness through temporal path-based accessibility for predefined groups, directly optimizing the dynamic activation of interactions over time to meet these fairness goals while minimizing resource usage. Moreover, in networked systems, fairness is a long-standing concern in resource allocation. This includes fair bandwidth sharing in communication networks [20], fair scheduling in wireless systems [16, 34], and equitable service provisioning. In edge-cloud environments, fair scheduling of updates (e.g., for recommender systems [17]) or computation offloading is critical. FML contributes to this line of work by providing a formal model for achieving group-level temporal reachability fairness specifically through the lens of minimizing discrete edge activations, a key proxy for communication cost, energy, or computational overhead, aligning with sustainability principles [25, 31]. Unlike other problems that focus on flow rates or latency targets, FML provides hard guarantees on temporal path existence for specified quotas from diverse groups.
@@ 4 The fair minimum labeling problem
The Fair Minimum Labeling (FML) problem addresses this gap by integrating three key aspects: (i) cost-aware edge activation planning, (ii) per-group temporal reachability guarantees, and (iii) flexible, group-specific coverage requirements. The goal is to compute a temporal edge labeling that minimizes total activation cost while satisfying fairness constraints across all groups (see Appendix˜B for an application example.)
@@
The FML problem generalizes both the Minimum Labeling (ML) [2] and the Minimum Steiner Labeling (MSL) [2] problems (see Observation˜1 in the appendix), which immediately implies that the FML problem is NP-complete. Moreover, we show that FML remains NP-hard and hard to approximate within a factor better than $\Omega(\log n)$ even in the special case where the terminal set is a singleton, i.e., $\mathcal{T}=\{t\}$.
@@
The result follows from a gap-preserving reduction from the Set Cover problem.
@@ 5 Approximation
Given the hardness results in Theorem˜1, we now present approximation algorithms that are tight in expectation with respect to cost and solve the FML problem in instances with a single terminal. We focus on the case where each node has color $B$, $R$, or no color, i.e., $\mathcal{C}=\{B,R,\emptyset\}$, and use set notation $B\subseteq V$ and $R\subseteq V$ to directly refer to the nodes of each group of colored nodes. Furthermore, we define the group-specific requirement functions as $f_{B}(B)=\alpha|B|$ and $f_{R}(R)=\alpha|R|$ for a fixed parameter $\alpha\in\mathbb{R}_{\geq 0}$, indicating that an equal fraction $\alpha$ of both $B$ and $R$ nodes must reach the terminal. This two-group setting captures a common fairness use case and provides a natural foundation for future extensions to multiple groups or terminals.
@@
Input: Graph $G=(V,E)$, terminal $B\subseteq V$, $R\subseteq V$, $t\in V$, $\alpha\in\mathbb{R}$
@@
Algorithm˜1 outlines our general framework for approximating the Fair Minimum Labeling (FML) problem using probabilistic tree embeddings. We begin with the input graph $G=(V,E)$ and the FML parameters ($R,B,t,\alpha$). We define the standard shortest path metric $d_{G}$ on $G$, where $d_{G}(u,v)$ is the minimum number of edges on a path between $u$ and $v$. We then employ the probabilistic tree embedding algorithm of Fakcharoenphol, Rao, and Talwar [13]. In practice, we first sample a single weighted tree $T=(V,E_{T},w)$ from the distribution $\mathcal{D}$ and then solve an adapted version of the FML problem on the sampled weighted tree $T$. We use the original node set $V$ with colors $B$ and $R$, and the parameters $\alpha$ and $t$. The objective, however, is modified to align with the structure provided by the embedding and the goal of minimizing the final cost in $G$. Specifically, we aim to find a temporal labeling $\lambda_{T}$ for the tree $T$ that minimizes the total weighted cost
$C_{T}=\sum_{(\{u,v\},t)\in\lambda_{T}}w(u,v)=\sum_{(\{u,v\},t)\in\lambda_{T}}d_{T}(u,v)$
subject to the coverage constraints being satisfied with respect to temporal reachability within the tree $T$. Specifically, fractions $\alpha|B|$ and $\alpha|R|$ of the nodes $B$ and $R$ can temporally reach the terminal $t$.
@@
  • An exact algorithm (Section˜5.1) finds the minimum weighted cost $C_{T,opt}$ satisfying the constraints exactly on $T$.
  • An approximation algorithm (Section˜5.2) finds a labeling $\lambda_{T}$ with weighted cost $C_{T}\leq C_{T,opt}$ that satisfies the constraints approximately, introducing an error factor $\xi\leq(1+\varepsilon)^{H+1}$ on the counts $b$ and $r$, where $H$ is the height of $T$ and $\varepsilon>0$ is the approximation parameter.
@@ 5.1 An exact algorithm for FML in weighted trees
We compute labels at the nodes of the tree $T$ in a bottom up fashion starting from the leafs. The edges of the tree have weights $w(u,v)\geq 1$ corresponding the shortest paths distance between $u,v\in V$ in $G$. Without loss of generality the maximum distance (in terms of edge weights) from any leaf to the root is at most $k$. Furthermore, all leafs are colored blue or red. We use labels of the form $(b,r,c)$ where $b$ and $r$ are the numbers of blue and red nodes in the subtree, and $c$ is the sum of edge weights used to connect the $b+r$ nodes to the node having the label. For example, a blue node gets a label $(1,0,0)$ and a red node $(0,1,0)$. Let $v$ be an inner node with $\ell$ children $u_{1},\ldots,u_{\ell}$. We construct the following labels based on the labels of the children. Let $\{L_{1},\dots,L_{\ell}\}$ be the sets of labels at $u_{1},\ldots,u_{\ell}$. We only choose at most one label per child that determines the (possibly empty) subtree to include. Therefore, we construct at most $\prod_{i=1}^{j=\ell}(1+|L_{j}|)$ labels at node $v$. At each node we keep at most one label $(b,r,c)$ for each possible pair $(b,r)$ with the minimum value of $c$. Because $r$ and $b$ are at most $n$ there are $\mathcal{O}(n^{2})$ possibilities of $(b,r)$ which is also an upper bound of the number of (non-dominated) labels at each node. Algorithm˜2 in the appendix shows the pseudocode.

### v2

@@ Abstract
Balancing resource efficiency and fairness is critical in networked systems that support modern learning applications. We introduce the Fair Minimum Labeling (FML) problem: the task of designing a minimum-cost temporal edge activation plan that ensures each group of nodes in a network has sufficient access to a designated target set, according to specified coverage requirements. FML captures key trade-offs in systems where edge activations incur resource costs and equitable access is essential, such as distributed data collection, update dissemination in edge-cloud systems, and fair service restoration in critical infrastructure. We show that FML is NP-hard and $\Omega(\log|V|)$-hard to approximate, where $V$ is the set of nodes, and we present probabilistic approximation algorithms that match this bound, achieving the best possible guarantee for the activation cost. We demonstrate the practical utility of FML in a fair multi-source data aggregation task for training a shared model. Empirical results show that FML enforces group-level fairness with substantially lower activation cost than baseline heuristics, underscoring its potential for building resource-efficient, equitable temporal reachability in learning-integrated networks.
@@ 1 Introduction
The introduction of FML as a novel optimization problem is motivated by both practical relevance and theoretical gaps. First, it applies to a wide range of systems where temporal resource allocation directly affects group equity. In distributed data gathering for learning applications, such as those involving environmental sensors, mobile agents, or geographically diverse data streams [31, 43, 42], FML enables minimal-cost activation plans that ensure all groups are fairly represented in the collected data. In edge-cloud systems, it helps decide when communication links should be active so that all device groups can receive timely model updates under bandwidth constraints. Similarly, in infrastructure restoration, FML can guide activation sequences that ensure fair reactivation of services. Second, FML fills a critical gap in the literature. Existing fairness-aware graph methods, such as fair clustering [24], node embeddings [37], and influence maximization [40, 44], are primarily designed for static settings and do not model temporal interactions or activation cost. Fairness has been studied in wireless systems [17], and temporal connectivity has been explored in dynamic networks [22, 27], but fairness constraints remain absent in algorithmic temporal graph design. FML addresses this gap by explicitly minimizing temporal edge activations, treated as a proxy for resource cost, while enforcing group-level temporal reachability. In contrast to decentralized approaches such as classical federated learning [19], where participation is uncoordinated, FML assumes centralized control over link activations, allowing the enforcement of fairness guarantees over time. This design objective distinguishes FML from prior work focused on static structures or continuous flows.
@@ 2 Preliminaries
A static graph is an ordered pair $G=(V,E)$, where $V$ is a finite set of nodes (or vertices) and $E\subseteq\{\{u,v\}\mid u,v\in V,u\neq v\}$ is a finite set of undirected edges. We assume static graphs are simple (no self-loops, no multiple edges between the same pair of nodes). A colored static graph $G=(V,E,c)$ with colors $\mathcal{C}$ has an additional function $c:V\rightarrow\mathcal{P}(\mathcal{C})$ assigning a (possibly empty) set of colors $c(u)$ to each node $u\in V$ where $\mathcal{C}=[c_{\max}]$ is a set of $c_{\max}$ colors. We use $V_{c}=\{v\in V\mid c(v)=c\}$. Let $n=|V|$ and m=|E|m=|{E}|.
A temporal graph is an ordered pair $\mathcal{G}=(V,\mathcal{E})$, where $V$ is a finite set of nodes and $\mathcal{E}$ is a finite set of temporal edges. Each temporal edge is a pair $e=(\{u,v\},\tau)$, where $\{u,v\}$ is an undirected pair of distinct nodes from $V$, and $\tau\in\mathbb{N}$ is the timestamp indicating when the interaction between $u$ and $v$ occurs.
Given a static graph $G=(V,E)$, a temporal labeling is a function $\lambda:E\rightarrow\mathcal{P}(\mathbb{N})$ that assigns a set of timestamps (possibly empty) to each static edge $e\in E$. A labeling $\lambda$ induces a temporal graph $G_{\lambda}=(V,E_{\lambda})$, where the node set is the same as the static graph’s, and the set of temporal edges is defined as: $E_{\lambda}=\{(\{u,v\},\tau)\mid e=\{u,v\}\in E\text{ and }\tau\in\lambda(e)\}$. The size (or cost) of a labeling $\lambda$, denoted by $|\lambda|$, is the total number of assigned timestamps, which corresponds to the number of temporal edges in the induced temporal graph, that is, $|\lambda|=|E_{\lambda}|=\sum_{e\in E}|\lambda(e)|$.
@@
By convention, we consider every node to be temporally reachable from itself (via a path of length zero), so $r(v,v)=1$ for all $v\in V$. The specific temporal graph $\mathcal{G}$ (often an induced graph $G_{\lambda}$) for which reachability is considered will always be clear from the context. Finally, for $U,W\subseteq V$, let $r(U,W)=\sum_{u\in U,v\in W}r(u,v)$ be the sum of pairwise reachability of nodes in $U$ to nodes in $W$.
@@ 3 Related Work
Temporal Graph Connectivity and Labeling. A core component of FML is finding a minimum-cost set of temporal edge activations. Mertzios et al. [27] (and its journal version  [28]) introduced the study of temporal network design in multi-labeled temporal graphs, where the goal is to optimize some global cost measure of the graph labeling, such as the total number of timestamps, while satisfying specific temporal reachability properties. The Minimum Labeling (ML) problem is such a temporal network design problem, seeking the minimum number of labels to make an entire static graph temporally connected. For its directed version on strongly connected digraphs, [27] had provided bounds on the number of labels needed, proving that 2 labels per arc and at most 4(|V|1)4(|V|-1) total labels are always sufficient. Using similar constructions on undirected graphs, Akrida et al. [2] proved that 2|V|32|V|-3 total labels are sufficient, later refined by Klobas et al. [25] to 2|V|42|V|-4 under certain conditions, who also showed ML is polynomially solvable. While ML aims for global temporal connectivity, FML focuses on targeted reachability for specific node groups to a designated terminal set, under explicit fairness (coverage) constraints for each group, and a cost minimization objective. Klobas et al. [25] also introduced extensions like Minimum Steiner Labeling (MSL), where the goal is to pairwise temporally connect a predefined set of terminal nodes. Our FML generalizes ML and MSL. Other research has explored concepts like temporally-connected spanning subgraphs [3], temporal core decompositions [15, 33], and various notions of temporal paths and reachability [22, 7, 29, 16, 34, 35, 30]. However, these works generally do not incorporate group fairness constraints or the specific cost model of minimizing temporal edge activations as FML does.
@@
Fairness in Graph Mining and Network Optimization. There is a significant body of work on incorporating fairness into static graph algorithms and network optimization, driven by the need for equitable outcomes in various applications. This includes, e.g., fair clustering [1, 10, 24], fair node embeddings [37, 6], fair influence maximization [40, 44], and graph learning tasks [9, 11, 32, 45, 36]. While these contributions are crucial for advancing fairness in graph-based AI, they predominantly operate on static graphs and often define fairness differently (e.g., proportionality in output, balanced representation in a static structure). For example, a related problem in static graphs is the Balanced Connected Subgraph (BCS) problem [5] which seeks a connected subgraph with balanced representation from different node colors. FML, in contrast, addresses fairness through temporal path-based accessibility for predefined groups, directly optimizing the dynamic activation of interactions over time to meet these fairness goals while minimizing resource usage. Moreover, in networked systems, fairness is a long-standing concern in resource allocation. This includes fair bandwidth sharing in communication networks [21], fair scheduling in wireless systems [17, 41], and equitable service provisioning. In edge-cloud environments, fair scheduling of updates (e.g., for recommender systems [18]) or computation offloading is critical. FML contributes to this line of work by providing a formal model for achieving group-level temporal reachability fairness specifically through the lens of minimizing discrete edge activations, a key proxy for communication cost, energy, or computational overhead, aligning with sustainability principles [26, 38]. Unlike other problems that focus on flow rates or latency targets, FML provides hard guarantees on temporal path existence for specified quotas from diverse groups.
@@ 4 The fair minimum labeling problem
The Fair Minimum Labeling (FML) problem addresses this gap by integrating three key aspects: (i) cost-aware edge activation planning, (ii) per-group temporal reachability guarantees, and (iii) flexible, group-specific coverage requirements. The goal is to compute a temporal edge labeling that minimizes total activation cost while satisfying fairness constraints across all groups (see Appendix˜B for an application example).
@@
The FML problem generalizes both the Minimum Labeling (ML) and the Minimum Steiner Labeling (MSL) problems (see Observation˜1 in the appendix). Since MSL is NP-complete [25], it immediately follows that FML is NP-complete. Moreover, we show that FML remains NP-hard and hard to approximate within a factor better than $\Omega(\log n)$ even in the special case where the terminal set is a singleton, i.e., $\mathcal{T}=\{t\}$.
@@ 5 Approximation
Given the hardness results in Theorem˜1, we now present approximation algorithms that are tight in expectation with respect to cost and solve the FML problem in instances with a single terminal. We focus on the case where each node has colors $B$ (blue), $R$ (red), or no color, i.e., $\mathcal{C}=\{B,R,\emptyset\}$. For convenience, we use the set notation $B\subseteq V$ and $R\subseteq V$ to directly refer to the nodes of each group of colored nodes. Furthermore, we define the group-specific requirement functions as $f_{B}(B)=\alpha|B|$ and $f_{R}(R)=\alpha|R|$ for a fixed parameter $\alpha\in\mathbb{R}$ with $0\leq\alpha\leq 1$, indicating that an equal fraction $\alpha$ of both $B$ and $R$ nodes must reach the terminal. This two-group setting captures a common fairness use case and provides a natural foundation for future extensions to multiple groups or terminals.
@@
Input: Graph $G=(V,E)$, $B\subseteq V$, $R\subseteq V$, terminal $t\in V$, $\alpha\in[0,1]$
@@
Algorithm˜1 outlines our general framework for approximating the Fair Minimum Labeling (FML) problem using probabilistic tree embeddings. We begin with the input graph $G=(V,E)$ and the FML parameters ($R,B,t,\alpha$). We define the standard shortest path metric $d_{G}$ on $G$, where $d_{G}(u,v)$ is the minimum number of edges on a path between $u$ and $v$. We then employ the probabilistic tree embedding algorithm of Fakcharoenphol, Rao, and Talwar [13]. In practice, we first sample a single weighted tree $T=(V,E_{T},w)$ from the distribution $\mathcal{D}$ and then solve an adapted version of the FML problem on the sampled weighted tree $T$. We use the original node set $V$ with colors $B$ and $R$, terminal $t$, and the parameter $\alpha$. The objective, however, is modified to align with the structure provided by the embedding and the goal of minimizing the final cost in $G$. We aim to find a temporal labeling $\lambda_{T}$ for the tree $T$ that minimizes the total weighted cost
$C_{T}=\sum_{(\{u,v\},\tau)\in\lambda_{T}}w(u,v)=\sum_{(\{u,v\},\tau)\in\lambda_{T}}d_{T}(u,v)$
subject to the coverage constraints being satisfied with respect to temporal reachability within the tree $T$, requiring that an $\alpha$-fraction of the nodes in each of B and R can temporally reach the terminal $t$.
@@
  • An exact algorithm (Section˜5.1) finds the minimum weighted cost $C_{T,\text{opt}}$ satisfying the constraints exactly on $T$.
  • An approximation algorithm (Section˜5.2) finds a labeling $\lambda_{T}$ with weighted cost $C_{T}\leq C_{T,\text{opt}}$ that satisfies the constraints approximately, introducing an error factor $\xi\leq(1+\varepsilon)^{H+1}$ on the counts of blue and red nodes, where $H$ is the height of $T$ and $\varepsilon>0$.
@@ 5.1 An exact algorithm for FML in weighted trees
We compute labels at the nodes of the tree $T$ in a bottom-up fashion starting from the leafs. The edges of the tree have weights $w(u,v)\geq 1$ corresponding the shortest paths distance between $u,v\in V$ in $G$. Without loss of generality the maximum distance (in terms of edge weights) from any leaf to the root is at most $k$. Furthermore, all leafs are colored blue or red. We use labels of the form $(b,r,c)$ where $b$ and $r$ are the numbers of blue and red nodes in the subtree, and $c$ is the sum of edge weights used to connect the $b+r$ nodes to the node having the label. For example, a blue node gets a label $(1,0,0)$ and a red node $(0,1,0)$. Let $v$ be an inner node with $\ell$ children $u_{1},\ldots,u_{\ell}$. We construct the following labels based on the labels of the children. Let $\{L_{1},\dots,L_{\ell}\}$ be the sets of labels at $u_{1},\ldots,u_{\ell}$. We only choose at most one label per child that determines the (possibly empty) subtree to include. Therefore, we construct at most $\prod_{i=1}^{j=\ell}(1+|L_{j}|)$ labels at node $v$. At each node we keep at most one label $(b,r,c)$ for each possible pair $(b,r)$ with the minimum value of $c$. Because $r$ and $b$ are at most $n$ there are $\mathcal{O}(n^{2})$ possibilities of $(b,r)$ which is also an upper bound of the number of (non-dominated) labels at each node. Algorithm˜2 in the appendix shows the pseudocode.